what is the diameter of a circle that has an arc with the measure of 100 degrees and an arc length of 55 pi cm?
Found 2 solutions by jedidiah, rajagopalan:
Answer by jedidiah(6)

Circumference= 2Pi D
Circumference= (360 degrees/ 100 degrees) * 55Pi cm = 198Pi cm
D= 198Pi/2Pi = 99 cm

Answer by rajagopalan(174)

Arc length=(100/360)*2*pi*R
(100/360)*2*pi*R = 50*pi
R=(360*50*pi)/(100*2*pi)
R=90
D=2*R=2*90=180
D=180 cm